How many moles of KCI would be produced from the decomposition of 4.39 moles of KClO 3 _ KClO 3 (s) _ KCl(s)+_ O 2 (g)
1
Expert's answer
2020-04-22T09:18:52-0400

Solution:

Balanced equation for the given chemical reaction:

2KClO3(s) --> 2KCl(s)+ 3O2(g)

According to the chemical equation: n(KClO3)/2 = n(KCl)/2 = n(O2)/3.

n(KCl) = n(KClO3) = 4.39 moles.

n(KCl) = 4.39 moles.


Answer: 4.39 moles of KCl would be produced from the decomposition of 4.39 moles of KClO3.